Indicate which of the following sets of lines are parallel, perpendicular, or neither.
y=2x-5
y=2x+5

3x+y=5
-x+3y=6

3x+y=5
x+y=3

Answer by edjones(8007)

1) parallel because they have the same slope, 2.
.
2)
3x+y=5
-x+3y=6
.
y=-3x+5
3y=x+6, y=x/3 + 2
.
They are perpendicular because the slope of one is the negative reciprocal of the other.
.
3)
3x+y=5
x+y=3
.
y=-3x+5
y=-x+3
.
Neither.
.
Ed